OpenStudy (starryknight):

Solve for y in the following system of equations. 6x - 4y = -12 -6x + 2y = 0 -6 2 6 4 can someone please break it down for me i dont understand this ( i take math online so i dont have anyone to explain this to me) free metal for best response)

OpenStudy (anonymous):

The solution for y is 6.

OpenStudy (starryknight):

thank you

OpenStudy (anonymous):

You find the solutions of BOTH equations and multiply them.

OpenStudy (anonymous):

Since y is when x = 0: 6x - 4y = -12 6(0) - 4y = -12 0 - 4y = -12 -4y = -12 y = 3 -6x + 2y = 0 -6(0) + 2y = 0 0 + 2y = 0 2y = 0 y = 2 Since the solutions to both are 3 and 2. Multiply them to get 3 x 2 = 6.